Two identical balls each have a mass of 10g. What charge should these balls be given so that their interaction equalises the force of universal gravitation acting between them? The radii of the balls may be ignored in comparision to distance between them?
 

Grade:12

1 Answers

Arun
5 years ago
Dear Anand
 
To exactly cancel the gravitational pull by electrostatic repulsion =
 
G m1 *m2 /r² = k * q²/r²
 
Now
 
Put the value of G and m1 and m2 and k.
 
q = 8.6 * 10^-13 C